# CompAgent: An Agentic Framework for Visual Compliance Verification

### v1

@@ Abstract
Visual compliance verification is a critical yet underexplored problem in computer vision, especially in domains such as media, entertainment, and advertising where content must adhere to complex and evolving policy rules. Existing methods often rely on task-specific deep learning models trained on manually labeled datasets, which are costly to build and limited in generalizability. While recent multi-modal large language models (MLLMs) offer broad real-world knowledge and policy understanding, they struggle to reason over fine-grained visual details and apply structured compliance rules effectively on their own. In this paper, we propose CompAgent, the first agentic framework for visual compliance verification. CompAgent augments MLLMs with a suite of visual tools—such as object detectors, face analyzers, NSFW detectors, and captioning models—and introduces a planning agent that dynamically selects appropriate tools based on the compliance policy. A verification agent then integrates image, tool outputs, and policy context to perform multi-modal reasoning. Experiments on public benchmarks show that CompAgent outperforms specialized classifiers, direct MLLM prompting, and curated routing baselines, achieving up to 7676% F1 score and a 1010% improvement over the state-of-the-art on the UnsafeBench dataset. Our results demonstrate the effectiveness of agentic planning and tool-augmented reasoning for scalable, accurate, and adaptable visual compliance verification.
@@ 1 Introduction
Existing approaches generally fall into two categories. The first relies on training dedicated deep learning classifiers to identify violations for predefined categories (e.g., nudity, violence, or hate symbols) [9, 36]. The second utilizes multi-modal large language models (MLLMs), either by crafting compliance-oriented prompts or by fine-tuning them for specific compliance rules [8, 24, 3, 12]. While these methods have shown promise, they face key limitations. Training domain-specific models is costly and impractical due to (a) the lack of labeled datasets for highly specialized compliance scenarios and (b) the rapid evolution of compliance policies, which renders static models obsolete. On the other hand, directly applying MLLMs often struggles to achieve optimal performance because (a) rules are often intricate and interdependent, making it challenging for a model to interpret them holistically, and (b) understanding fine-grained visual details—such as contextual cues or nuanced cultural indicators—requires structured guidance beyond naive prompting. Consequently, current solutions either incur prohibitive labeling costs or deliver suboptimal accuracy.
@@
  • Agentic Compliance Framework: We design the first ever (as per our knowledge at the time this work is done) agent-based system for visual compliance verification that dynamically orchestrates specialized vision tools (e.g., face detection, Not Safe for Work (NSFW) detection, object detection, image captioning, and Interactive Advertising Bureau (IAB) categorization) according to compliance rules.
  • Policy-Aware Agent: We develop an agent that parses compliance policies to identify applicable rules, plans and selects the most relevant tools, eliminating the need for manual mapping and finally performs verification using the image, policy, and tool outputs to jointly reason about compliance, improving both precision and interpretability.
@@
We evaluate CompAgent on two public datasets under diverse compliance policies and demonstrate that it significantly outperforms baselines and state-of-the-art methods, including direct MLLM prompting and specialized classifiers. Comprehensive ablation studies validate the effectiveness of our design choices, and qualitative analyses illustrate the explainability of our framework. Furthermore, although our current work focuses on single-image compliance, the architecture can seamlessly extend to video compliance verification with minimal modification, underscoring its generalizability.
@@ 2 Related Works
Figure 2: The architecture of CompAgent. Given compliance policies and an input image, it leverages toolbox-provided tools to extract metadata, reasons over the policy, and iteratively refines its process until a final decision is made.
Early image content moderation relied heavily on rule-based systems and traditional machine learning classifiers. These methods often focused on predefined heuristics, statistical models, or textual cues associated with images rather than complex visual analysis. The approaches evolved from simple keyword filtering and metadata analysis using rule-based systems [13] to more sophisticated methods using Machine Learning. In this section, we examine the related works in image content moderation techniques using machine learning models through two primary approaches: traditional computer vision frameworks and state-of-the-art Large Language Models (LLMs).
@@ 2.1 Traditional Computer Vision Frameworks
Historically, methods such as Support Vector Machines (SVMs) and Random Forests were used with handcrafted features to classify images as per content preferences. These systems relied on features such as color histograms, texture patterns, and edge detection to categorize content. Recent works using traditional computer vision often employ Convolutional Neural Networks (CNNs) for analyzing visual features, and learning their pattern with predefined safety policies, such as nudity, violence, or hate speech. Tools for unsafe image classification, such as NudeNet [22] and NSFW-Nets [7] are built on CNN architectures. Leveraging CNNs, several Contrastive Language–Image Pre-training (CLIP) based [28] approaches have been developed for image content moderation, such as [21, 29]. Notably, Poppi et al. [26] introduced HySAC, which innovatively uses CLIP embeddings in hyperbolic space to model content safety hierarchies, while maintaining awareness of unsafe concepts and enabling flexible moderation. Similarly, authors in [25] proposed SafeCLIP, which fine-tunes a CLIP backbone on synthetic safe/unsafe image-text pairs to “unlearn” links between NSFW concepts and the embedding space. In the realm of specialized detection, CNN-based approaches such as CUE-Net [30] have shown particular effectiveness in video violence detection, combining spatial cropping with transformer architectures to achieve state-of-the-art accuracy on standard benchmarks like RWF-2000 [6] and Real Life Violence Situations Dataset [31]. More comprehensive frameworks have also emerged, such as ICM-Assistant [33], which implements a rule-based system for transparent content filtering with structured outputs including classifications and explanations.
@@ 2.2 Large Language Model based Frameworks
LLM-based approaches have significantly influenced unsafe image detection by integrating multi-modal capabilities, enabling more nuanced and context-aware moderation compared to traditional computer vision methods. These models, often extended into vision-language models (VLMs), combine visual understanding with natural language processing to assess images for violations such as nudity, violence, or hate content, providing not just binary classifications but also detailed rationales and policy-compliant evaluations. [16] introduced LLaVA, one of the first large vision-language models with instruction-tuning capabilities, which became a crucial baseline for subsequent safety-focused research. LlavaGuard [11] introduces an open VLM-based framework for assessing image safety, offering customizable taxonomies, safety ratings, categories, and rationales to evaluate compliance with diverse policies, outperforming traditional safeguards in accuracy and flexibility. Authors in  [10] proposed ECSO, a training-free defense mechanism that leverages the inherent safety mechanisms of text-only LLM backbones by transforming images into textual descriptions when unsafe content is detected, showing substantial improvements in harmless rates. [34] introduces SafeVision, an image moderation system using vision-LLMs that combines human-like understanding with guardrails for harmful content detection. Research in  [17] involves the use of cross-modal co-embeddings in LLMs, and zero-shot learning. Another avenue of research focuses on supervised fine-tuning to adapt LLMs for specific content moderation tasks, and authors in  [18] delve into the pitfalls of fine-tuning for content moderation. In this work, we explore agentic LLM-based approaches for the corresponding application, and to the best of our knowledge, there does not exist significant research works focusing image content moderation from this perspective.
@@ 3 Method
We illustrate the proposed CompAgent framework in Fig. 2. The framework consists of a planning agent that orchestrates compliance verification through interactive tool-use. For any incoming image, it systematically analyzes compliance against predefined policy categories through a two-stage process: (1) tool selection and invocation based on policy requirements, and (2) compliance verification through multi-modal reasoning. The tool selection and verification procedures are summarized in Algorithm 1.
At the core of CompAgent is an LLM that acts as the planning agent, equipped with a diverse set of visual analysis tools ranging from basic content detection to specialized compliance verification models. When presented with an image and compliance policies, the agent first analyzes the policy requirements to determine necessary verification steps. It then strategically selects and invokes appropriate tools from its suite to gather evidence, maintaining an explicit reasoning chain throughout the process. The final compliance decision is made by integrating the tool outputs, image content, and policy context through structured multi-modal reasoning.
@@ 3.1 Tool Suite
CompAgent leverages a carefully curated set of specialized tools for comprehensive content analysis, focusing on efficiency and accuracy rather than exhaustive tool application. Unlike existing approaches that employ large, potentially redundant tool collections, our framework adopts a minimal but sufficient tool set organized into three categories:
@@
Summarization Tools. The foundation of our analysis pipeline consists of tools that provide high-level content understanding. Image Summarization tool [1] is to generate natural language descriptions capturing key visual elements and their relationships. It processes each image through multiple visual models to generate comprehensive descriptions that capture both obvious and subtle content elements, serving as the primary tool for establishing overall context and identifying areas that require deeper analysis.
@@
Content Detection Tools. This category comprises fundamental analysis tools that extract specific content elements. Face Detection [2] processes facial attributes while maintaining privacy, extracting metadata such as age ranges, expressions, and emotional attributes without identifying individuals. Object Detection [2] performs comprehensive scene analysis, identifying and localizing objects within images. It returns structured lists of detected objects along with their confidence scores, supporting analysis of prohibited items or dangerous materials. Text Detection [2] employs Optical Character Recognition (OCR) capabilities to extract and process textual content from images, performing word-level detection. This enables analysis of text-based violations such as hate speech or inappropriate messages. Content Moderation [1] identifies potentially unsafe content categories. It processes images through multiple moderation models and returns structured labels indicating the presence and severity of policy-relevant content.
@@
Specialized Compliance Tools. These tools are specifically designed for compliance verification tasks. LlavaGuard Classification [11] functions as an expert system for compliance verification, providing comprehensive assessments including safety ratings (Safe/Unsafe), violation categories, and detailed rationales. It operates as a specialized system trained specifically on compliance policies and verification tasks. SafeCLIP [37] Analysis performs zero-shot detection of toxic content across seven predefined categories including explicit content, violence, and prohibited substances. It is particularly reliable for initial content screening. ICM Assistant [33] employs template-based verification to provide structured, explainable safety assessments. It matches visual content against predefined violation templates and generates detailed explanations linking specific visual elements to policy violations, supporting transparent and auditable verification decisions.
Each tool outputs standardized formats containing detection results, confidence measures, and structured metadata, facilitating seamless integration of evidence during final compliance assessment. Rather than applying every tool to each image, CompAgent selectively invokes specific tools based on policy requirements and initial content analysis, ensuring thorough yet efficient verification.
3.2 Planning Agent
CompAgent orchestrates the compliance verification process through a ReAct (Reasoning and Acting) [35] framework using carefully designed prompts that encode complete policy specifications and structured instructions for systematic analysis. When presented with an image and compliance policies, the agent systematically analyzes policy requirements, strategically selects and invokes appropriate tools, integrates their outputs, and maintains an explicit reasoning chain throughout the verification process. The agent’s ability to decompose complex compliance requirements into specific verification tasks, combined with its strategic tool selection and systematic reasoning process, enables robust and adaptable content moderation across diverse use cases. After gathering evidence through selected tools, the agent integrates multiple information streams: the original image content, tool outputs, the complete compliance policy specification, and its own reasoning chain to produce a structured assessment. This assessment includes a binary safety rating (Safe/Unsafe), identification of any violated policy categories, and a detailed rationale explaining the decision. Through this combination of strategic planning, tool orchestration, and systematic reasoning, CompAgent achieves effective and explainable compliance verification while minimizing computational overhead, making it suitable for real-world applications where both accuracy and performance are critical.
@@ 3.2 Planning Agent
5:function AgentVerification($s_{0}$)
6:  while not terminal state do
7:   $s_{t}\leftarrow$ CurrentState() $\triangleright$ Current reasoning state
8:   $a_{t}\leftarrow$ SelectTool($s_{t},T$) $\triangleright$ Select next tool
9:   if $a_{t}$ = CONCLUDE then
10:     return MakeSafetyAssessment($s_{t}$)
11:   end if
12:   $s_{t+1}\leftarrow$ UpdateState($s_{t}$, ExecuteTool($a_{t},I$))
13:  end while
14:end function
15:function SelectAction($s_{t},T$)
16:  return Agent($s_{t}$).decide() $\triangleright$ Agent decides to use tool or conclude
17:end function
18:function MakeSafetyAssessment($s_{t}$)
19:  return {ClassifySafety($s_{t}$), DetermineCategory($s_{t}$), GenerateRationale($s_{t}$)}
20:end function
The implementation of CompAgent leverages LangGraph [14] utilizing a ReAct-style prompt template that enables iterative reasoning and tool use. Each tool is implemented as a function that interfaces with underlying models and returns structured results. The agent employs MLLM as its core language model, with a system prompt that guides the iterative process of analyzing requirements, selecting tools, and making final compliance decisions. The implementation follows a modular design where tool invocations, reasoning steps, and final assessments are tracked and logged systematically, enabling both real-time monitoring and post-hoc analysis of the verification process.
The CompAgent framework offers several distinct advantages over traditional compliance verification methods. The tool-augmented approach achieves improved accuracy through specialized tool selection tailored to specific compliance requirements, while providing enhanced interpretability through explicit reasoning chains that document the verification process. A crucial feature is its training-free operation — the system requires no labeled data or fine-tuning, making it particularly well-suited for real-world deployment scenarios where compliance policies frequently evolve. Through system prompts that include compliance requirements, the framework maintains flexibility to adapt to new policies while ensuring consistent evaluation across diverse content types. This combination of strategic planning, tool orchestration, and systematic reasoning enables CompAgent to achieve effective and explainable compliance verification while minimizing computational overhead through targeted tool invocation rather than applying all tools indiscriminately.

### v2

@@ Abstract
Visual compliance verification is a critical yet underexplored problem in computer vision, especially in domains such as media, entertainment, and advertising where content must adhere to complex and evolving policy rules. Existing methods often rely on task-specific deep learning models trained on manually labeled datasets, which are costly to build and limited in generalizability. While recent Multimodal Large Language Models (MLLMs) offer broad real-world knowledge and policy understanding, they struggle to reason over fine-grained visual details and apply structured compliance rules effectively on their own. In this paper, we propose CompAgent, the first agentic framework for visual compliance verification. CompAgent augments MLLMs with a suite of visual tools—such as object detectors, face analyzers, NSFW detectors, and captioning models—and introduces a planning agent that dynamically selects appropriate tools based on the compliance policy. A compliance verification agent then integrates image, tool outputs, and policy context to perform multimodal reasoning. Experiments on public benchmarks show that CompAgent outperforms specialized classifiers, direct MLLM prompting, and curated routing baselines, achieving up to 7676% F1 score and a 1010% improvement over the state-of-the-art on the UnsafeBench dataset. Our results demonstrate the effectiveness of agentic planning and robust tool-augmented reasoning for scalable, accurate, and adaptable visual compliance verification.
Warning: This paper contains imagery and content that some readers may find sensitive or disturbing.
@@ 1 Introduction
Existing approaches generally fall into two categories. The first relies on training dedicated deep learning classifiers to identify violations for predefined categories (e.g., nudity, violence, or hate symbols) [9, 40]. The second utilizes multimodal large language models (MLLMs), either by crafting compliance-oriented prompts or by fine-tuning them for specific compliance rules [8, 26, 3, 13]. While these methods have shown promise, they face key limitations. Training domain-specific models is costly and impractical due to (a) the lack of labeled datasets for highly specialized compliance scenarios and (b) the rapid evolution of compliance policies, which renders static models obsolete. On the other hand, directly applying MLLMs often struggles to achieve optimal performance because (a) rules are often intricate and interdependent, making it challenging for a model to interpret them holistically, and (b) understanding fine-grained visual details—such as contextual cues or nuanced cultural indicators—requires structured guidance beyond naive prompting. Consequently, current solutions either incur prohibitive labeling costs or deliver suboptimal accuracy.
@@
  • Agentic Compliance Framework: We design the first ever (as per our knowledge at the time this work is done) agent-based system for visual compliance verification that dynamically orchestrates specialized vision tools (e.g., face detection, object detection, and more) according to compliance rules.
  • Policy-Aware Agents: We develop agents that parse compliance policies to identify applicable rules, plan and select the most relevant tools, eliminating the need for manual mapping and finally perform verification using the image, policy, and tool outputs to jointly reason about compliance, improving both precision and interpretability.
@@
We evaluate CompAgent on two public datasets under diverse compliance policies and demonstrate that it significantly outperforms baselines and state-of-the-art methods, including direct MLLM prompting and specialized classifiers. Comprehensive ablation studies validate the effectiveness of our design choices, and qualitative analyses illustrate the explainability of our framework. Although our current work focuses on single-image verification, the modular and agentic design of CompAgent lays a practical foundation for future extensions to video compliance verification, highlighting its strong potential to generalize to video applications.
@@ 2 Related Works
Figure 2: The architecture of CompAgent. Given compliance policies and an input image, it leverages provided tools to extract metadata, reasons over the policy, and iteratively refines its process until a final decision is made.
Early image content moderation relied heavily on rule-based systems and traditional machine learning classifiers. These methods often focus on predefined heuristics, statistical models, or textual cues associated with images rather than complex visual analysis. The approaches evolve from simple keyword filtering and metadata analysis using rule-based systems [14] to more sophisticated methods using Machine Learning. In this section, we examine the related works in image content moderation techniques using machine learning models through two primary approaches: traditional computer vision frameworks and state-of-the-art Large Language Models (LLMs).
@@ 2.1 Traditional Computer Vision Frameworks
Historically, methods such as Support Vector Machines (SVMs) and Random Forests were used with handcrafted features to classify images as per content preferences. These systems relied on features such as color histograms, texture patterns, and edge detection to categorize content. Recent works using traditional computer vision often employ Convolutional Neural Networks (CNNs) for analyzing visual features, and learning their pattern with predefined safety policies, such as nudity, violence, or hate speech. Tools for unsafe image classification, such as NudeNet [24] and NSFW-Nets [7] are built on CNN architectures. Leveraging CNNs, several Contrastive Language–Image Pre-training (CLIP) based [30] approaches have been developed for image content moderation, such as [23, 31]. Notably,  [28] introduces HySAC, which innovatively uses CLIP embeddings in hyperbolic space to model content safety hierarchies, while maintaining awareness of unsafe concepts and enabling flexible moderation. Similarly, authors in [27] propose Safe-CLIP, which fine-tunes a CLIP backbone on synthetic safe/unsafe image-text pairs to “unlearn” links between NSFW concepts and the embedding space. In the realm of specialized detection, CNN-based approaches such as CUE-Net [32] have shown particular effectiveness in video violence detection, combining spatial cropping with transformer architectures to achieve state-of-the-art accuracy on standard benchmarks like RWF-2000 [6] and Real Life Violence Situations Dataset [35]. More comprehensive frameworks have also emerged, such as ICM Assistant [37], which implements a rule-based system for transparent content filtering with structured outputs.
@@ 2.2 Large Language Model based Frameworks
LLM-based approaches have significantly influenced unsafe image detection by integrating multimodal capabilities, enabling more nuanced and context-aware moderation compared to traditional computer vision methods. These models, often extend into vision-language models (VLMs), combine visual understanding with natural language processing to assess images for violations such as nudity, violence, or hate content, providing not just binary classifications but also detailed rationales and policy-compliant evaluations. [17] introduces LLaVA, one of the first large VLMs with instruction-tuning capabilities, which becomes a crucial baseline for subsequent safety-focused research. LlavaGuard [12] introduces an open VLM-based framework for assessing image safety, offering customizable taxonomies, safety ratings, categories, and rationales to evaluate compliance with diverse policies, outperforming traditional safeguards in accuracy and flexibility. Authors in  [10] propose ECSO, a training-free defense mechanism that leverages the inherent safety mechanisms of text-only LLM backbones by transforming images into textual descriptions when unsafe content is detected, showing substantial improvements in harmless rates. [38] introduces SafeVision, an image moderation system using vision-LLMs that combines human-like understanding with guardrails for harmful content detection. Research in [18] involves the use of cross-modal co-embeddings in LLMs, and zero-shot learning. Another avenue of research focuses on supervised fine-tuning to adapt LLMs for specific content moderation tasks, and authors in [19] delve into the pitfalls of fine-tuning for content moderation.
2.3 Agentic AI based Solutions
Recent advances in LLMs have catalyzed the emergence of agentic systems across diverse application domains. In spatial reasoning and 3D understanding, VADAR [20] demonstrates how agentic frameworks can dynamically generate APIs to decompose complex visual queries into executable programs, achieving superior performance on benchmarks requiring multi-step grounding and inference. Similarly, in video understanding, Agent-of-Thoughts Distillation [33] employs specialized agents for temporal grounding, object detection, and action recognition to construct reasoning chains for VideoQA tasks. The information retrieval domain has witnessed the rise of Agentic Retrieval-Augmented Generation (Agentic RAG) [34], where autonomous agents dynamically manage retrieval strategies and iteratively refine contextual understanding through reflection, planning, and tool use. More recently, Agentic Deep Research systems [41] integrate reasoning agents with search capabilities to tackle complex, multi-step information synthesis tasks. Even in structured database interactions, agentic workflows have proven effective, with divide-and-conquer strategies and verification steps enhancing Text-to-SQL generation [11].
Despite this proliferation of agentic approaches, the application of autonomous agents to visual compliance verification remains entirely unexplored. To the best of our knowledge, we are the first to introduce an agentic framework specifically designed for automated visual compliance verification, addressing the unique challenges of decomposing compliance criteria, orchestrating evidence gathering, and providing interpretable verification traces.
@@ 3 Method
We illustrate the proposed CompAgent framework in Fig. 2. The framework consists of three main components: (1) Planning Agent that interprets policy requirements from text descriptions and the input image, then strategically selects appropriate visual tools from (2) Tool Suite containing basic detectors and specialized compliance models to gather evidence through interactive invocation, and finally (3) Compliance Verification Agent (CVAgent) that integrates the collected tool outputs with image content and policy context through structured multimodal reasoning to produce the final compliance judgment. The overall procedure is summarized in Algorithm 1.
3.1 Planning Agent
The Planning Agent serves as the strategic orchestrator of the compliance verification process, implemented through a ReAct (Reasoning and Acting) [39] framework with carefully designed prompts that encode complete policy specifications and structured instructions for systematic analysis. When presented with an image and compliance policies, the Planning Agent maintains the current reasoning state $s_{t}$ at time $t$ and systematically analyzes policy requirements to determine the next action $a_{t}$.
At each step, the Planning Agent operates on the current state $s_{t}=\{I,P,E_{t}\}$, which contains the input image $I$, compliance policy $P$, and accumulated evidence $E_{t}$ from previously executed tools. Based on this state and the available Tool Suite $T$, the agent strategically selects the next tool to invoke or decides to conclude the verification process. The agent updates the state after each tool execution: $s_{t+1}\leftarrow$ UpdateState($s_{t}$, ExecuteTool($a_{t},I$)), integrating new evidence into the reasoning chain. When the Planning Agent determines that sufficient evidence has been gathered to make an informed compliance decision, it selects the terminal action $a_{t}=$ CONCLUDE, triggering the transition to the verification phase handled by CVAgent.
The agent’s decision-making follows an explicit reasoning chain that decomposes complex compliance requirements into specific verification tasks, identifying which tools from the Tool Suite are most appropriate for gathering relevant evidence. The Planning Agent’s tool selection strategy is guided by multiple factors: the policy requirements that need verification, the agent’s understanding of each tool’s capabilities and limitations through comprehensive tool descriptions, the evidence already gathered from previous tool invocations, and the current context of the analysis. Rather than applying every tool exhaustively, the Planning Agent selectively invokes specific tools based on necessity, ensuring thorough yet efficient verification. For instance, if the policy concerns age-restricted content, the agent may prioritize face detection tools; if text-based violations are relevant, Text Detection becomes critical.
@@ 3.2 Tool Suite
The Tool Suite $T$ comprises a carefully curated collection of specialized tools that the Planning Agent can invoke during the compliance verification process. Each tool in the Tool Suite is defined by two key components: (1) a tool description that includes comprehensive information about its input arguments, return format, usage guidelines, and its specific capabilities, optimal usage scenarios, and known limitations—these detailed descriptions are critical for enabling the Planning Agent’s intelligent tool selection strategy described in the previous section, and (2) an invoke function that executes the tool on the input image and returns structured results. When the Planning Agent selects a tool $a_{t}$ from the Tool Suite, it calls ExecuteTool($a_{t},I$) which triggers the tool’s invoke function with the input image $I$, producing outputs in a standardized format containing detection results, confidence measures, and structured metadata. This standardization facilitates seamless integration of evidence during state updates and final compliance assessment.
@@
The Tool Suite is organized into three categories, focusing on efficiency and accuracy rather than exhaustive tool application. Summarization Tools [1] form the foundation of our analysis pipeline, providing high-level content understanding and serving as the primary mechanism for establishing overall context. They leverage multiple visual models to generate comprehensive natural language descriptions that capture both obvious and subtle content elements.
@@
Content Detection Tools. This category comprises fundamental analysis tools that extract specific content elements. Face Detection [2] processes facial attributes while maintaining privacy. Object Detection [2] identifies and localizes objects within images with confidence scores. Text Detection [2] employs Optical Character Recognition capabilities to extract and process textual content from images, performing word-level detection. Content Moderation [1] identifies potentially unsafe content categories. It processes images through multiple moderation models and returns structured labels indicating the presence and severity of policy-relevant content.
@@
Specialized Compliance Tools. These tools are specifically designed for compliance verification tasks. LlavaGuard Classification [12] functions as an expert system for compliance verification, providing comprehensive assessments including safety ratings, violation categories, and detailed rationales. It operates as a specialized system trained specifically on compliance policies and verification tasks. Safe-CLIP [42] performs zero-shot detection of toxic content across seven predefined categories including explicit content, violence, and prohibited substances. ICM Assistant [37] employs template-based verification to provide structured, explainable safety assessments. It matches visual content against predefined violation templates and generates detailed explanations linking specific visual elements to policy violations, supporting transparent and auditable verification decisions.
3.3 Compliance Verification Agent
The CVAgent performs the final compliance assessment after the Planning Agent concludes evidence gathering. Implemented as an MLLM, the CVAgent receives the final state $s_{t}=\{I,P,E_{t}\}$ containing the input image $I$, compliance policy $P$, and all accumulated evidence $E_{t}$ from executed tools. Through step-by-step reasoning over the state, image, and policy, the CVAgent integrates these multiple information streams through structured multimodal reasoning to produce a comprehensive safety assessment $A=\{rating,category,rationale\}$, where ratingrating is a binary classification (Safe/Unsafe), categorycategory identifies the specific policy violation if any, and rationalerationale provides a detailed explanation of the decision.
The MLLM-based CVAgent synthesizes evidence by analyzing the image content, examining tool outputs with their confidence measures, interpreting the policy specification, and tracing the Planning Agent’s reasoning chain in $E_{t}$. This step-by-step analysis enables cross-validation of findings, conflict resolution, and evidence weighting based on tool confidence and policy relevance. The agent maintains transparency by generating detailed rationales that explicitly link visual evidence to policy requirements, explaining which content elements triggered violations and why.
@@ 3.3 Compliance Verification Agent
5:function CompAgent($s_{0}$)
6:  $t=0$
7:  while not terminal state do
8:   $s_{t}\leftarrow$ CurrentState() $\triangleright$ Current reasoning state
9:   $a_{t}\leftarrow$ PlanningAgent($s_{t},T$) $\triangleright$ Select next tool
10:   if $a_{t}$ = CONCLUDE then
11:     return CVAgent($s_{t}$) $\triangleright$ Verification agent assesses
12:   end if
13:   $s_{t+1}\leftarrow$ UpdateState($s_{t}$, ExecuteTool($a_{t},I$))
14:   $t\leftarrow t+1$
15:  end while
16:end function
17:function PlanningAgent($s_{t},T$)
18:  return Agent($s_{t}$).decide() $\triangleright$ Planning agent decides to use tool or conclude
19:end function
20:function CVAgent($s_{t}$)
21:  return {ClassifySafety($s_{t}$), DetermineCategory($s_{t}$), GenerateRationale($s_{t}$)}
22:end function
The implementation of CompAgent leverages LangGraph [15] with a ReAct-style prompt template that enables iterative reasoning and tool use. The Planning Agent can be implemented using either an LLM or MLLM as its core language model, while the CVAgent is implemented as an MLLM to enable direct image analysis. The Planning Agent’s prompt encodes complete policy specifications and tool descriptions, enabling systematic analysis and strategic tool selection, while the CVAgent’s prompt structures the final assessment process to integrate evidence, image content, and policy context.